Joe traveled against the wind in a small plane for 3 hr. The return trip with the wind took 2.8 hr. Find the speed of the wind if the speed of the plane in still air is 180mph.
Answer by venugopalramana(3286)

Joe traveled against the wind in a small plane for 3 hr. The return trip with the wind took 2.8 hr. Find the speed of the wind if the speed of the plane in still air is 180mph.
let the wind speed be = x mph
speed of the plane in still air is 180mph...
hence speed against the wind =180-x
time of travel against wind =3 hrs
we have distance travelled=d =speed*time=3*(180-x)=540-3x
speed pro wind =180+x
time of travel pro wind=2.8 hrs
hence distance travelled=d =speed*time=2.8*(180+x)=504+2.8x
so 504+2.8x=540-3x
2.8x+3x=540-504=36
5.8x=36
x=36/5.8=360/58=180/29 mph